Please help me to solve the following;
The complement of an angle measures 10(degrees) less than one-fifth of its supplement. Find the measure of the angle.
I started out by writing: 90-x = 10 -1/5(90 - x)
I"m not sure how to handle the fraction (one-fifth). The answer is x=80. Everything I've tried did not work. Your help is appreciated.
Thanks
Joshua

Answer by mananth(16946)

angle =x
complement= (90-x)
supplement = 180-x
(90-x)= 1/5(180-x)-10
multiply the equation by 5
5(90-x)= (180-x)-50
450-5x=180-x-50
5x-x= 450-180+50
4x=320
/4
x=80 deg